# OceanGPT: A Large Language Model for Ocean Science Tasks

### v6

@@ 5.1 Insights from Performance Results
Figure 5: Ocean task-level results. Left: Automatic evaluation. Right: Human evaluation. Compared to baselines, OceanGPT performs better than llama2-chat-7b, vicuna-1.5-7b and chatglm2-6b in both two settings. The instance-level result is in Table 11 (Appendix A.2).
@@
In Figure 5, we compare the performance of OceanGPT with the three baseline models across 15 sub-tasks at the task-level in the ocean domain. We utilize both automatic and human evaluators, then compute the win rate (%) with baseline models. Compared to the baselines (llama2-chat-7b, vicuna-1.5-7b, chatglm2-6b) [^5] We have trained OceanGPT-7B, thus we only compare open-sourced LLMs around 7B. We are running resources to train OceanGPT-30B, and will compare with more LLMs in the future., OceanGPT outperforms in the majority of tasks, which demonstrates the effectiveness of the proposed approach.
Figure 6: Evaluation results of OceanGPT in the ocean science tasks in OceanBench. The complete experimental results are shown in the Table 10 (Appendix A.2).
@@
As shown in Figure 7, we design three indicators to measure the data generation effect of our proposed method from the perspectives of knowledge quality, expertise and diversity. We use manual evaluation to calculate the scores of the three indicators from 1 to 5. The higher the score, the better the effect of the testing model. It can be seen that the evolving generator agent can effectively enhance the richness of ocean data. When the extraction agent is at work, the expertise of the content is greatly improved. At the same time, the inspector agent plays a significant role in enhancing the quality of the generated data. The results show that multi-agent collaboration is effective for ocean instruction generation.
@@ 5.2 Exploring the Potential of OceanGPT
In this section, we explore the potential of OceanGPT from the perspectives of ocean science and ocean engineering. For ocean science (Section 5.2.1), we focus on the key scientific issues of nuclear pollution in the ocean environment. For ocean engineering (Section 5.2.2), we use the Gazebo  [^6] https://github.com/uuvsimulator/uuv_simulator as the simulator to test OceanGPT’s ability to control underwater robots.
@@ 5.2.1 OceanGPT for Ocean Science
In Figure 8, given the question “How to conduct research on interfacial chemistry and toxicological effects of key radioactive nuclides”, we compare the outputs of our model and vicuna-1.5-7b. The result shows that OceanGPT demonstrates a higher level of knowledge expertise when describing the content of radioactive nuclide research. Its textual content is not only clear in structure and well-organized, but also covers various aspects of radioactive nuclide research, from experimental design to data analysis, and then to risk assessment and disposal guidelines. In contrast, although vicuna-1.5-7b has clear expression and logicality, it lacks depth and specific content related to radioactive nuclides. Overall, OceanGPT has advantages in terms of knowledge expertise, quality, and richness. The complete outputs of both models are shown in the Table 4.
Figure 8: Case analysis on ocean science task. We use blue font to represent the different texts by OceanGPT. The instruction is: How to conduct research on interfacial chemistry and toxicological effects of key radioactive nuclides?
Figure 9: Our model can be instructed for underwater robot control in the simulation platform of Gazebo, which shows OceanGPT gains preliminary embodied intelligence capabilities.
@@ 5.2.2 OceanGPT for Ocean Engineering
Ocean engineering focuses on the design, development, and management of structures and systems within the ocean environment. It plays an indispensable role in harnessing the vast potential of the oceans while ensuring sustainable and secure maritime operations. To facilitate interaction between OceanGPT and the external world, we integrate machine code instructions into its training data. As depicted in Figure 9, OceanGPT can instruct underwater robots via code or console commands, allowing them to execute basic path-finding operations. In this example, by using natural language as a prompt, the OceanGPT can automatically generate code (the robot generate a double helix path) for underwater robot to operate based on human instructions. In fact, the experimental result suggests that OceanGPT has the potential to acquire embodied intelligence. Though we make preliminary attempts for ocean robot interaction, it paves the way for future advanced oceanic models to undertake intricate robotic control and planning tasks.
@@ 6 Conclusion
In this paper, we introduce OceanGPT, the first-ever oceanographic pre-trained language model, which is expert in various ocean science tasks. To alleviate the difficulties for obtaining ocean data, we propose an domain construction framework called DoInstruct, which constructs the ocean instruction dataset by multi-agent collaboration. Each agent in our designed framework is considered as an expert in a specific topic and is responsible for generating the corresponding data. Our generated dataset consists of diverse instructions to align the desired behaviors in ocean science issues. Additionally, we establish the first oceanography benchmark, OceanBench, to evaluate the capabilities of LLMs in ocean domain. Though comprehensive analysis, we observe that OceanGPT not only demonstrates a higher level of knowledge expertise for oceans science tasks but also gains preliminary embodied intelligence capabilities in ocean engineering. We will continue to improve OceanGPT by training on larger corpus with larger models (e.g., 30B, 70B) and maintain OceanBench by adding new data and tasks.
Ethical Considerations
The datasets and model in this paper are designed for exploratory analysis. Within the domain of LLMs, the distribution of both pre-training data and instruction data might exhibit biases, potentially influencing the model’s outputs. Throughout experiments, we have carefully examined the sampled data, eliminating any content that could be deemed toxic or offensive. Yet, it’s possible that OceanGPT might occasionally yield incorrect or imprecise results. Hence, users leveraging the insights provided by this model should exercise their discretion and bear the risk.
Reproducibility Statement
Pre-training and instruction-tuning: https://github.com/zjunlp/KnowLM.
Instruction processing: https://github.com/zjunlp/EasyInstruct (The code is still under organization and will be integrated into EasyInstruct soon.).
Model will be released at HuggingFace soon.
@@ Appendix A Appendix
A.1 Limitations and Ethical Concerns
A.1.1 Bias in Data Distribution
A.1.2 Hallucination in Large Language Models
A.2 Complete Results in OceanBench
Figure 11: Instance-level results (automatic evaluation) in 15 tasks of OceanBench.
Task Num Task Num Task Num
Analysis 674 Classification 895 Judgment 655
Commonsense Reasoning 1,024 Open-ended Generation 930 Extraction 1,078
Recommendation 1,089 Description 1,246 Summary 149
Editing 1,075 Identification 464 Transformation 401
Question Answering 1,230 Letter Writing 359 Others 157
Table 1: The statistics of OceanBench. Num denotes the number of samples.
Figure 12: Distribution of our OceanBench.
A.3 Implementation Details
Hyperparameter
@@ A.3 Implementation Details
Prompt for ”Fine-Tuned Agent as the Literature Extractor”:
@@
Prompt for ”Evolving Agent as the Generator”:
@@
Please check if following sentences contain rich ocean related information. If so, output ”related”. Otherwise, output ”unrelated”.
@@
The study of interfacial chemistry and toxicological effects of key radioactive nuclides is a complex research field that requires a multidisciplinary approach. It involves multiple techniques such as experimental research, theoretical research, and numerical simulations. The following is a detailed outline of the steps to conduct research on the interface chemistry and toxicological effects of key radioactive nuclides: 1. Determine the specific radioactive nuclides of interest: Different radioactive nuclides have different chemical properties and toxicological effects. Therefore, it is necessary to select specific radioactive nuclides of interest and establish a research plan. 2. Conduct laboratory experiments: Laboratory experiments are crucial for studying the chemical properties and toxicological effects of radioactive nuclides. The experiments involve measuring the chemical properties of radioactive nuclides, such as their reactivity, stability, and chemical reactivity. 3. Develop computational models: Computational models can be used to simulate the chemical and toxicological effects of radioactive nuclides. This involves developing mathematical models and algorithms to predict the behavior of radioactive nuclides under different conditions. 4. Analyze experimental data: Experimental data obtained from laboratory experiments need to be analyzed and interpreted. This involves studying the experimental results and comparing them with theoretical predictions. 5. Develop toxicological risk assessment methods: Toxicological risk assessment methods are used to evaluate the potential hazards of radioactive nuclides. This involves developing models that can predict the toxicological effects of radioactive nuclides and assessing their risks to human health. 6. Conduct field experiments: Field experiments are necessary to study the behavior of radioactive nuclides in real-world environments. This involves measuring the chemical and toxicological effects of radioactive nuclides in different ecosystems and studying their distribution and transport. 7. Develop guidelines for safe disposal: Safe disposal of radioactive nuclides is critical for preventing environmental contamination and protecting human health. This involves developing guidelines for the safe disposal of radioactive nuclides and their storage and transport.
vicuna-1.5-7b
@@
Conducting research on interfacial chemistry and toxicological effects of key radioactive nuclides requires a multidisciplinary approach, combining expertise in chemistry, physics, biology, and environmental science. Here are some general steps you can follow to conduct research in this area: 1. Identify the key radioactive nuclides of interest: Start by identifying the specific radioactive nuclides that you want to study. This may include elements such as uranium, plutonium, and cesium, among others. 2. Develop a research question: Once you have identified the key radioactive nuclides, develop a research question that addresses an important aspect of their interfacial chemistry and toxicological effects. This question should be specific and focused, and should guide your research efforts. 3. Design experiments: Develop a set of experiments that will help you answer your research question. This may involve conducting laboratory experiments, field studies, or both. Be sure to choose experiments that are appropriate for your research question and that use methods that are reliable and valid. 4. Collect and analyze data: Once you have conducted your experiments, collect and analyze the data you have generated. Be sure to carefully document your methods and results, and to interpret your data in a clear and concise manner. 5. Interpret results and draw conclusions: Use your data to draw conclusions about the interfacial chemistry and toxicological effects of the key radioactive nuclides. Be sure to consider any limitations of your study and to consider how your findings may be generalized to other contexts. 6.Communicate your results: Finally, share your findings with others through publication in academic journals, presentations at conferences, and other means. Be sure to clearly communicate the importance of your work and its implications for interfacial chemistry and toxicology.

### v7

@@ 5.1 Insights from Performance Results
In Figure 5, we compare the performance of OceanGPT with the three baseline models across 15 sub-tasks at the task-level in the ocean domain. We utilize both automatic and human evaluators, then compute the win rate (%) with baseline models. Compared to the baselines (llama2-chat-7b, vicuna-1.5-7b, chatglm2-6b), OceanGPT outperforms in the majority of tasks, which demonstrates the effectiveness of the proposed approach.
Figure 6: Evaluation results of OceanGPT in the ocean science tasks in OceanBench. The complete experimental results are shown in Figure 12 (Appendix A).
Figure 7: Performance analysis for different agents. We design three indicators to measure the generation effect.
@@
As shown in Figure 7, we design three indicators to measure the data generation effect of our proposed method from the perspectives of knowledge quality, expertise and diversity. We use manual evaluation to calculate the scores of the three indicators from 1 to 5. The higher the score, the better the effect of the testing model. It can be seen that the evolving generator agent can effectively enhance the richness of ocean data. When the extraction agent is at work, the expertise of the content is greatly improved. At the same time, the inspector agent plays a significant role in enhancing the quality of the generated data. It shows that multi-agent collaboration is effective for ocean instruction generation.
@@ 5.2 Exploring the Potential of OceanGPT
In this section, we explore the potential of OceanGPT from the perspectives of ocean science and ocean engineering. For ocean science (Section 5.2.1), we focus on the key scientific issues of nuclear pollution in the ocean environment. For ocean engineering (Section 5.2.2), we explore the potential in robotics applications (Li et al., 2023). Specifically, we use Gazebo  [^4] https://github.com/uuvsimulator/uuv_simulator as the simulator (Manhães et al., 2016) to test OceanGPT’s ability to control underwater robots.
Figure 8: Case analysis on ocean science task. We use blue font to represent the difference and the instruction is: How to conduct research on interfacial chemistry and toxicological effects of key radioactive nuclides?
Figure 9: Our model can be instructed for underwater robot control in the simulation platform of Gazebo which shows OceanGPT gains preliminary embodied intelligence capabilities.
@@ 5.2.1 OceanGPT for Ocean Science
In Figure 8, we compare the outputs of OceanGPT and vicuna-1.5-7b. It shows that OceanGPT shows a higher level of knowledge expertise when describing the content of radioactive nuclide research. Its textual content is not only clear in structure and well-organized, but also covers various aspects of radioactive nuclide research, from experimental design to data analysis, and then to risk assessment and disposal guidelines. In contrast, although vicuna-1.5-7b has clear expression and logicality, it lacks depth and specific content related to radioactive nuclides. Overall, OceanGPT has advantages in terms of knowledge expertise, quality, and richness. The complete outputs are shown in the Table 6.
@@ 5.2.2 OceanGPT for Ocean Engineering
Ocean engineering focuses on the design, development, and management of structures and systems within the ocean environment. It plays an indispensable role in harnessing the vast potential of the oceans while ensuring sustainable and secure maritime operations. To facilitate interaction between OceanGPT and the external world, we synthesize robotic code data and integrate those machine code instructions into the training data.
As depicted in Figure 9, OceanGPT can instruct underwater robots via code or console commands, allowing them to execute basic path-finding operations. In this example, by using programming language as a prompt, our OceanGPT can automatically generate code (the robot generate a double helix path) for underwater robot to operate complex tasks (based on human instructions). In fact, the experimental result suggests that OceanGPT has the potential to acquire embodied intelligence. Though we make preliminary attempts for ocean robot interaction, it paves the way for advanced oceanic models to undertake intricate robotic control and complex planning tasks.
@@ 6 Conclusion
In this paper, we introduce OceanGPT, the first-ever oceanographic pre-trained language model, which is expert in various ocean science tasks. To alleviate the difficulties for obtaining ocean data, we propose an domain construction framework called DoInstruct, which constructs the ocean instruction dataset by multi-agent collaboration. Each agent in our designed framework is considered as an expert in a specific topic and is responsible for generating the corresponding data. Our generated dataset consists of diverse instructions to align the desired behaviors in ocean science issues. Though comprehensive analysis, we observe that OceanGPT not only demonstrates a higher level of knowledge expertise for oceans science tasks but also gains preliminary embodied intelligence capabilities in ocean engineering. We will continue to improve OceanGPT by training on larger corpus with larger models (e.g., 30B, 70B) and maintain OceanBench by adding new data and tasks.
Limitations
Bias in Data Distribution
Hallucination in LLMs
Acknowledgments
We would like to express gratitude to the anonymous reviewers for constructive comments. This work was supported by the National Natural Science Foundation of China (No. 62206246), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), CCF-Baidu Open Fund, the “Pioneer and Leader + X” Plan Project of Zhejiang Province under Grant (2024C01162), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
@@ Appendix A Appendix
Hyperparameter Setting
@@
Figure 10: Instance-level results (automatic evaluation)
Figure 11: Distribution of our OceanBench.
The Cost for Fine-tuning GPT-3.5-Turbo
For fine-tuning GPT-3.5-turbo, we use the reference code provided by OpenAI to fine-tune their language model. Overall, during the actual debugging process, we train and test the model multiple times, spending a total of nearly 500 US dollars (with the number of high-quality training samples being around 2000). Each time we run the script to train the model, it takes several hours.
The training cost is 0.008 USD per 1K tokens, the input cost during use is 0.012 USD per 1K tokens, and the output cost is 0.016 USD per 1K tokens. Assuming our prompt’s input and output for one conversation is 1000 tokens, and if we have 2000 training samples with actual testing on 10000 samples, our training cost would be approximately 16.8 USD. The usage cost of the model after fine-tuning is about 138.0 USD, making the total cost around 154.8 USD. Since we debugged multiple times in the actual process, the real expenditure is greater. Overall, the overall training cost is not high and is affordable.
Comparison between Our Fine-tuning Method and the Prefix Prompts
In the paper, we define 5 marine science topics, but this is a very broad categorization. In reality, each major topic contains many subtopics. For example, the topic ’Ecology and Environment’ includes subtopics like marine meteorology, marine pollution, and over a dozen others. Altogether, these subtopics amount to over 500. Each of these subtopics is relatively independent and very important. Concatenating them as a prefix to GPT-3.5-turbo would exceed its maximum length limit and the actual usage cost would also be significant. Therefore, we believe that fine-tuning GPT-3.5-turbo is a better choice. The prompt examples are shown in Table 3 and Table 4.
Instruction: You are a helpful ocean assistant. You are to extract the question from the provided content.
Input: Raw sentences in the marine literature (The instruction prompt will be concatenated with raw sentences about seawater resources ).
Output: Answer: Existing methods of seawater resource exploitation have many problems, such as causing soil erosion and environmental pollution. Therefore, we need to seek more sustainable development methods, including water conservation, wastewater recycling, and the development of new water resources. Question: Please discuss your views on the current methods of developing seawater resources.
Table 3: The prompt for fine-tuning GPT-3.5-turbo.
Instruction: You are a helpful ocean assistant. You are to extract the question from the provided content.
Input: Raw sentences in the marine literature (The instruction prompt will be concatenated with raw sentences about seawater resources ).
The demonstration and answer pairs: I will first give you some typical examples to help you become a marine expert. Demonstration 1: … Answer 1: … Demonstration 2: … Answer 2: … Demonstration 3: … Answer 3: … Demonstration 4: … Answer 4: … (The demonstration and answer pairs for each marine subtopics. over 500 sub-categories. Each sub-categories has different task types )
Output: Answer: … Question: … (Concatenating them as a prefix to GPT-3.5-turbo would exceed its maximum length limit and the actual usage cost is significant )
Table 4: The prefix prompt to GPT-3.5-turbo.
The Similarity Calculating Method in the Deduplication Procedure
Because comparing pairs for similarity involves a significant number of calculations, we choose a simple and effective method to address this challenge. We primarily use hash detection to compare two samples. First, we pre-extract keywords from the question part of each sample and then combine them into a new string. For example, the keywords for a data sample might be ’advice’, ’ocean’, and ’nuclear leakage’. We then employ hash detection to compare the keywords of the two samples. This method can relatively accurately prevent data leakage during the training process. It’s important to note that sometimes the extraction of keywords can lead to redundancy or repetition, so we sometimes process them multiple times. Additionally, we also randomly select some samples and use the GPT-3.5-turbo API for detection to check for any cases of incomplete processing.
Additionally, regarding the deduplication process between the benchmark and our training dataset, we remove only a hundred or two hundred samples from the training set in the actual experiment, which is not a large number.
Prompt for "Fine-Tuned Agent as the Literature Extractor":
@@
Prompt for "Evolving Agent as the Generator":
@@
Please check if following sentences contain rich ocean related information. If so, output "related". Otherwise, output "unrelated".
@@
OceanGPT The study of interfacial chemistry and toxicological effects of key radioactive nuclides is a complex research field that requires a multidisciplinary approach. It involves multiple techniques such as experimental research, theoretical research, and numerical simulations. The following is a detailed outline of the steps to conduct research on the interface chemistry and toxicological effects of key radioactive nuclides: 1. Determine the specific radioactive nuclides of interest: Different radioactive nuclides have different chemical properties and toxicological effects. Therefore, it is necessary to select specific radioactive nuclides of interest and establish a research plan. 2. Conduct laboratory experiments: Laboratory experiments are crucial for studying the chemical properties and toxicological effects of radioactive nuclides. The experiments involve measuring the chemical properties of radioactive nuclides, such as their reactivity, stability, and chemical reactivity. 3. Develop computational models: Computational models can be used to simulate the chemical and toxicological effects of radioactive nuclides. This involves developing mathematical models and algorithms to predict the behavior of radioactive nuclides under different conditions. 4. Analyze experimental data: Experimental data obtained from laboratory experiments need to be analyzed and interpreted. This involves studying the experimental results and comparing them with theoretical predictions. 5. Develop toxicological risk assessment methods: Toxicological risk assessment methods are used to evaluate the potential hazards of radioactive nuclides. This involves developing models that can predict the toxicological effects of radioactive nuclides and assessing their risks to human health. 6. Conduct field experiments: Field experiments are necessary to study the behavior of radioactive nuclides in real-world environments. This involves measuring the chemical and toxicological effects of radioactive nuclides in different ecosystems and studying their distribution and transport. 7. Develop guidelines for safe disposal: Safe disposal of radioactive nuclides is critical for preventing environmental contamination and protecting human health. This involves developing guidelines for the safe disposal of radioactive nuclides and their storage and transport.
@@
vicuna-1.5-7b Conducting research on interfacial chemistry and toxicological effects of key radioactive nuclides requires a multidisciplinary approach, combining expertise in chemistry, physics, biology, and environmental science. Here are some general steps you can follow to conduct research in this area: 1. Identify the key radioactive nuclides of interest: Start by identifying the specific radioactive nuclides that you want to study. This may include elements such as uranium, plutonium, and cesium, among others. 2. Develop a research question: Once you have identified the key radioactive nuclides, develop a research question that addresses an important aspect of their interfacial chemistry and toxicological effects. This question should be specific and focused, and should guide your research efforts. 3. Design experiments: Develop a set of experiments that will help you answer your research question. This may involve conducting laboratory experiments, field studies, or both. Be sure to choose experiments that are appropriate for your research question and that use methods that are reliable and valid. 4. Collect and analyze data: Once you have conducted your experiments, collect and analyze the data you have generated. Be sure to carefully document your methods and results, and to interpret your data in a clear and concise manner. 5. Interpret results and draw conclusions: Use your data to draw conclusions about the interfacial chemistry and toxicological effects of the key radioactive nuclides. Be sure to consider any limitations of your study and to consider how your findings may be generalized to other contexts. 6.Communicate your results: Finally, share your findings with others through publication in academic journals, presentations at conferences, and other means. Be sure to clearly communicate the importance of your work and its implications for interfacial chemistry and toxicology.
@@
Task Type Example
Analysis Input: Analyze the impact of marine environmental dynamics processes on the wave environment. Output: The dynamic processes of the oceanic environment have a significant influence on the wave conditions, such as the temperature of seawater, salinity, and ocean current affecting the height and frequency of waves. Meanwhile, the wave conditions can further impact the dynamics of ocean current and tides.
Classification Input: Classify the ships in the East China Sea according to their type. Output: Ships in the East China Sea can be classified into categories such as merchant ships, fishing vessels, passenger ships, and warships, etc. Merchant ships are primarily used for cargo transport, fishing vessels for fishery resources, passenger ships for tourism and personnel transport, and warships for maritime security and defense. These diverse types of vessels are frequently active in the waters of the East China Sea.
Commonsense Reasoning Input: Why are fish resources in the East China Sea gradually decreasing? Output: Fish resources in the East China Sea are gradually decreasing, mainly due to the impact of factors such as overfishing and environmental pollution. With the increase in population and economic development, fishing activities have been continuously strengthened, resulting in over-exploitation of fish resources. In addition, marine environmental pollution also has adverse effects on the survival and reproduction of fish resources. Therefore, protecting the marine environment and rationally developing fishing resources are key to ensuring the sustainable development of fish resources in the East China Sea.
Description Input: Describe the classification and distribution areas of coral reefs. Output: Coral reefs are calcareous submarine structures formed by coral entities and their attachments. According to their crescent characteristics, they can be categorized into large ring coral reefs, small ring coral reefs, crisscross serrated coral reefs, irregular shaped coral reefs, etc. Approximately 70% of the world’s coral reefs are distributed in tropical and subtropical sea areas, with the most important distribution areas being the western Indian Ocean, western Pacific Ocean, and eastern Pacific Ocean.
Editing Input: Edit a big data report used for analyzing marine ecosystems. Output: Important parameters that can be tracked in marine ecosystems include water temperature, salinity, direction of waves, quantity and species of plankton, rate of ocean acidification, etc. In marine big data analysis, we use these data to train machine learning algorithms to understand the changes in the marine ecosystem, forecast potential environmental changes, etc. By utilizing these critical parameters, we can perceive the panorama and pattern of the marine ecosystem and extract valuable information from it to support adaptive and sustainable management practices in the ocean.
Extraction Input: Extract the definition of the greenhouse effect and its causes. Output: The greenhouse effect refers to a ẗhermal blanketf̈ormed by the absorption and reflection of solar heat radiation by greenhouse gases (such as carbon dioxide, methane, fluorocarbons, etc.) in the atmosphere, thereby raising the surface temperature of the earth. The sources of greenhouse gases include human activities, natural gases, deforestation, etc., which leads to an intensification of the greenhouse effect and global warming.
Judgment Input: Is active intervention suitable for maintaining the coral reef ecosystem? Output: Active intervention can protect coral reef ecosystems, such as through artificial restoration, controlling sea water pollution, controlling carbon emissions, and reasonable development of coral reefs. Therefore, the coral reef ecosystem is suitable for active intervention and maintenance.
Letter Writing Input: Write a letter to the local environmental protection agency about the waste discharge and pollution recently discovered in the sea area, proposing solutions. Output: Dear xx Environment Bureau: Recently, we have discovered a large amount of waste emission and pollution in the sea area, which has caused serious damage to our marine environment. We suggest that you take a series of measures to strengthen the monitoring and management of pollution sources in the sea area, punish illegal emissions in accordance with the law, and educate illegal polluting enterprises to enhance their environmental protection awareness. We hope that you can take action to effectively protect our marine environment.
Open-ended Generation Input: Briefly introduce the principle and method of wave energy generation. Output: The principle of wave power generation is to convert wave energy into mechanical energy, and then turn the mechanical energy into electrical energy through a generator. The methods of wave power generation include: floating system, wave influx system, wave reflux system, and wave pressure system.
Question Answering Input: What is three-dimensional digital ocean modeling technology? Output: Three-dimensional marine digital modeling technology involves the use of advanced remote sensing, lidar, sonar, and other detection technologies to acquire surface and subsurface oceanographic topographic data. Through digital modeling methods, a three-dimensional mathematical model of the ocean can be constructed. This type of model can provide information on topography, ocean currents, hydrology and acoustics, which is beneficial for applications in ocean resource development, environmental protection, and management, among others.
Recognition Input: Identify the following sea fish species: sea bream, codfish, saury, and shark. Output: Dorada is a common type of sea fish, laterally compressed in form with a big mouth, and is known for its delicious meat. Cod is also a common sea fish, slender in shape with a soft hanging lower jaw, and is also a delicacy. Saury is a type of fish that enters ports from summer to autumn and can be made into a variety of dishes. Sharks are carnivorous marine animals with many species, some of which are primarily used as food, while others may attack humans.
Recommendation Input: Recommend a desalination method suitable for deep-sea water resource development Output: The development of deep-sea water resources is suitable for the adoption of low-temperature seawater desalination technology, which uses cooling water to lower the seawater temperature below the ambient temperature, condensing out fresh water. Compared to traditional seawater desalination technology, low-temperature seawater desalination can conserve energy and reduce costs, while the resulting brine is also more concentrated.
Summary Input: Provide a summary of the current status and future development prospects of offshore wind power generation technology. Output: Ocean wind power generation technology is one of the hot technologies in the field of renewable energy, its power generation efficiency is becoming increasingly high, equipment is becoming increasingly advanced, and the costs are gradually decreasing. It has a broad development prospect and can inject new development momentum into China’s marine economy.
Transformation Input: How to transform marine plants into green products? Output: The conversion process of marine plants can be divided into initial collection and extraction of effective parts, followed by processing to obtain the final product. Among them, extraction and processing are critical steps, and they vary for different marine plants. Common processing methods include alkali extraction, acid extraction, and enzyme extraction. The resulting green products also vary, such as chlorophyll, peptide substances, polysaccharide substances, etc.
Table 7: Examples for tasks in OceanBench.